OpenStudy (anonymous):

During the winter, an ice cream shop has 45% of the employees it has during the summer. Which equation shows that during the summer, the ice cream shop has 20 employees?

OpenStudy (anonymous):

Summeremployees = 20 shame on the mathteacher who gace you that problem....

OpenStudy (anonymous):

gace = gave

OpenStudy (stormfire1):

I suspect there was a list of possible answers that wasn't provided here...

OpenStudy (anonymous):

20=0.45*s 45=0.20*s w=20%*45 0.45*20

OpenStudy (anonymous):

that is the list of possible answers

OpenStudy (anonymous):

please help smart please